A computer company receives 360 applications from computer graduates for a job planning a line of new Web servers. Suppose that 210 of these applicants majored in computer science, 147 majored in business and 51 majored in both subjects. How many of these applicants majored neither in computer science nor in business?
1
Expert's answer
2020-04-07T16:42:00-0400

Number of applicants majored only in computer science"=210-51=159"

Number of applicants majored only in business"=147-51=96"

Number of applicants in both"=51"

Total number of applicants"=360"

So number of applicants majored in neither computer science nor business"=360-210-147+51=54"